The length of a rectangle is 5in. more than twice its width. If the perimeter of the rectangle is 52 in., find the width of the rectangle.
Answer by checkley75(3666)

LENGTH=(2W+5)
2L+2W=52
2(2W+5)+2W=52
4W+10+2W=52
6W=52-10
6W=42
W=42/6
W=7 INCHES IS THE WIDTH.
PROOF
2(2*7+5)+2*7=52
2(14+5)+14=52
2*19+14=52
38+14=52
52=52